Circular No. 7966

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2002 Q2 (LINEAR) AND C/2002 Q3 (LINEAR) S. S. Sheppard and D. C. Jewitt, University of Hawaii, report that

CCD images of C/2002 Q3 on Sept. 4.3 and 5.3 UT taken with the 2.2-m reflector at Mauna Kea show three separate fragments a few arcsec apart; the brightest, component A (R = 21.7; photometry through a 1′′-radius aperture), appears elongated; component B was at R = 22.1 and component C at 22.6. The astrometric positions are given on MPEC 2002-R20.

Z. Sekanina, Jet Propulsion Laboratory, communicates: “Applying my model for the split comets and using a code developed recently by P. W. Chodas and myself, I determined from 46 differential astrometric observations of the C/2002 Q2-C/2002 Q3 pair, taken between Aug. 26 and Sept. 1 (MPECs 2002-Q37, 2002-Q50, 2002-Q51, 2002-R19, and 2002-R20), that C/2002 Q2 is the main comet, with companion C/2002 Q3 having broken off from it in 1969.2 ± 5.1 yr at a distance of some 57 ± 6 AU from the sun and > 30 AU south of the ecliptic. The separation velocity was only 0.15 ± 0.04 m/s, and the secondary comet has been subjected to a relative deceleration of 33.2 ± 5.1 units of 10^{−5} solar gravitational acceleration. The mean residual from the solution is ± 1′′.03. The time of separation may be even more uncertain than indicated, if the orbits depart significantly from parabolas. Because of too-limited information on the C/2002 Q3 system (see above) available at this time, the fragmentation sequence for its three components remains unclear, with two extreme scenarios possible: either they all separated from C/2002 Q2 a long time ago (say, before ∼ 1985) or two of them were released very recently (during the past several months) from the third. If so, component C is the parent to component B and perhaps to component A as well, although A could have alternatively separated from B after B broke off from C. Continuing observations of these components should soon constrain the range of possible scenarios. If the relative configuration of A, B, and C remains essentially unchanged over a period of a few weeks, they all are old fragments. If their separations begin to increase rapidly (by a factor of ∼ 2 or more in a couple of weeks) and the line joining them rotates toward the north-south direction, they are products of recent events. The predicted separation distances and position angles (for 0^{h} TT) of C/2002 C3 from C/2002 C2 are: Sept. 3, 3137′′, 304^{o}.5; 8, 3630′′, 314^{o}.0; 13, 4110′′, 321^{o}.3; 18, 4423′′, 326^{o}.5; 23, 4452′′, 330^{o}.0; 28, 4207′′, 332^{o}.2; Oct. 3, 3800′′, 333^{o}.7.”

2002 September 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green